OpenStudy (anonymous):

Help. Police can estimate the speed of a vehicle before the brakes are applied using the formula 0.75d=s^2/30.25. where s is the speed in miles per hour and D is the length of the vehicles skid marks. what was the approximate speed of a vehicle that left a skid of 150 feet. A)11 miles per hour B)15 Miles per hour C)58 miles per hour D)67 miles per hour I know its eitherC or D, I just am not sure how to find the answer! Thanks everyone!

OpenStudy (anonymous):

plug 150 in for d so .75(150)=s^2/30.25

OpenStudy (anonymous):

\[0.75\times 150=\frac{s^2}{30.25}\] solve for \(s\)

OpenStudy (anonymous):

112.5=s^2/30.45 Multiply both sides by 30.25 3,403.125=s^2

OpenStudy (anonymous):

s=the square root of 3403.125

OpenStudy (anonymous):

this equals 58.3363 mph

OpenStudy (anonymous):

I just got that as well! thank you so much for setting it up! it makes more sense now!